# ALPCAHUS: Subspace Clustering for Heteroscedastic Data

### v2

@@ Abstract
Principal component analysis (PCA) is a key tool in the field of data dimensionality reduction. Various methods have been proposed to extend PCA to the union of subspace (UoS) setting for clustering data that comes from multiple subspaces like $K$-Subspaces (KSS). However, some applications involve heterogeneous data that vary in quality due to noise characteristics associated with each data sample. Heteroscedastic methods aim to deal with such mixed data quality. This paper develops a heteroscedastic-focused subspace clustering method, named ALPCAHUS, that can estimate the sample-wise noise variances and use this information to improve the estimate of the subspace bases associated with the low-rank structure of the data. This clustering algorithm builds on K-Subspaces (KSS) principles by extending the recently proposed heteroscedastic PCA method, named LR-ALPCAH, for clusters with heteroscedastic noise in the UoS setting. Simulations and real-data experiments show the effectiveness of accounting for data heteroscedasticity compared to existing clustering algorithms. Code available at https://github.com/javiersc1/ALPCAHUS.
Index Terms:
Heteroscedastic data, heterogeneous data quality, subspace bases estimation, subspace clustering, union of subspace model, unsupervised learning.
@@ I Introduction
Many modern data science problems require learning an approximate signal subspace basis for some collection of data. This is important for downstream tasks involving subspace basis coefficients such as classification [1], regression [2], and compression [3]. Besides subspace learning, one may be interested in clustering data points that originate from multiple subspaces. Formally, subspace clustering, or union of subspace (UoS) modeling, is an unsupervised machine learning problem where the goal is to cluster unlabeled data and find the subspaces associated with each data cluster. When the cluster assignments are known, it is easy to find the subspaces, and vice versa. This problem becomes nontrivial when both components must be estimated [4]. This clustering problem has many applications, such as image segmentation [5], motion segmentation [6], image compression [7], and system identification [8].
@@
Some applications involve heterogeneous data samples that vary in quality due in part to noise characteristics associated with each sample. Some examples of heteroscedastic datasets include environmental air quality data [9], astronomical spectral data [10], and biological sequencing data [11]. In heteroscedastic settings, the noisier data samples can significantly corrupt the basis estimates [12]. In turn, this corruption can worsen clustering performance. Popular clustering methods such as Sparse Subspace Clustering (SSC) [13], $K$-Subspaces (KSS) [14], and Subspace Clustering via Thresholding (TSC) [15] implicitly assume that data quality is consistent. For example, in SSC, the method relies on the self-expressiveness property of data that requires using other samples to estimate similarity. From our experiments, we found that this implicit data quality assumption can degrade clustering quality for heteroscedastic data.
@@
Figure 1: Two 1D subspaces, colored blue and yellow, with data consisting of two noise groups shown with circle and triangle marker types.
@@
Because of these limitations, we developed a subspace clustering algorithm, inspired from KSS principles, that explicitly models noise variance terms, without assuming that data quality is known. The method adaptively clusters data while learning noise characteristics. See Fig. 1 for a visualization where Ensemble KSS (EKSS) [16] returns poor subspace bases estimates whereas our method found more accurate subspace bases and improved clustering quality.
@@
We extend our previous work [17] by generalizing the LR-ALPCAH formulation to the UoS setting for clustering heteroscedastic data. The proposed approach achieved ${\sim}3$ times lower clustering error than existing methods, and it achieved relatively low clustering error even when very few high quality samples are available. The paper is divided into a few key sections. Section II introduces the heteroscedastic problem formulation for subspace clustering. Section III discusses related work in subspace clustering and reviews the heteroscedastic subspace algorithm LR-ALPCAH. Section IV introduces the proposed subspace clustering method named ALPCAHUS. Section V covers synthetic and real data experiments that illustrate the effectiveness of modeling heteroscedasticity in a clustering context. Sec. VI discusses some limitations of our method and possible extensions.
@@ II Problem Formulation
Let $K$ denote the number of subspaces that is either known beforehand or estimated using other methods. Before describing the general union of subspace model, the single subspace model ($K=1$) is introduced.
II-A Single-Subspace Model ($K=1$)
Let $\bm{y}_{i}\in\mathbb{R}^{D}$ denote the data samples for index $i\in\{1,\ldots,N\}$, where $D$ denotes the ambient dimension. Let $\bm{x}_{i}$ represent the low-dimensional data sample generated by $\bm{x}_{i}=\bm{U}\bm{z}_{i}$ where $\bm{U}\in\mathbb{R}^{D\times d}$ is an unknown basis for a subspace of dimension $d$ and $\bm{z}_{i}\in\mathbb{R}^{d}$ are the corresponding basis coordinates. Then the heteroscedastic model we consider is
@@ II-A Single-Subspace Model ($K=1$)
assuming Gaussian noise with variance $\nu_{i}$, where $\bm{I}$ denotes the $D\times D$ identity matrix. We consider two cases: one where each data sample may have its own noise variance, and one where the case where there are $G$ groups of data having shared noise variance terms $\{\nu_{1},\ldots,\nu_{G}\}$. Sec. III-B discusses an optimization problem based on this model that estimates the heterogeneous noise variances $\{\nu_{i}\}$ and the subspace basis $\bm{U}$.
@@ II-B Union of Subspaces Model ($K\geq 1$)
Let $\bm{Y}=\left[\begin{matrix}\bm{y}_{1}&\ldots&\bm{y}_{N}\end{matrix}\right]\in\mathbb{R}^{D\times N}$ denote a matrix whose columns consist of all $N$ data points $\bm{y}_{i}\in\mathbb{R}^{D}$. We generalize (1) to model the data with a union of subspaces model as follows
@@
where $\bm{U}_{k}\in\mathbb{R}^{D\times d_{k}}$ is a subspace basis that has subspace dimension $d_{k}$. Here $\bm{z}_{i}\in\mathbb{R}^{d_{k}}$ denotes the basis coefficients associated with $\bm{x}_{i}$, and $\bm{\epsilon}_{i}\in\mathbb{R}^{D}$ denotes noise for that point drawn from $\bm{\epsilon}_{i}\sim\mathcal{N}(\bm{0},\nu_{i}\bm{I})$.
@@
$c_{i}=\operatorname*{arg\,min}_{k}\|\bm{y}_{i}-\bm{U}_{k}\bm{U}_{k}^{T}\bm{y}_{i}\|_{2}^{2},\quad\forall\bm{y}_{i}\in\bm{Y}.$ (3)
This label describes the subspace association of a data point $\bm{y}_{i}$ that has the lowest residual between the original sample and its reconstructed value given the subspace model. In general, the goal is to estimate all of the subspace bases $\mathcal{U}=(\bm{U}_{1},\ldots,\bm{U}_{K})$ and cluster assignments $\mathcal{C}=(c_{1},\ldots,c_{N})$ that describe the subspace labels of all points.
@@ III-A Subspace Clustering
Many subspace clustering algorithms fall into a general umbrella of categories such as algebraic methods [18], iterative methods [19], statistical methods [20], and spectral clustering methods [21] [22]. In recent years, both spectral clustering-based methods and iterative methods have become popular.
@@ III-A1 Spectral Clustering
Many methods build on spectral clustering. This method is often used for clustering nodes in graphs. Spectral clustering aims to find the minimum cost “cuts” in the graph to partition nodes into clusters. Given some collection of data points, one way to construct a graph is to assume that “nearby” data samples are highly connected in the graph. Thus, it is possible to construct a graph from data samples, with various levels of connectedness, by using some metric to assign affinity/similarity for each pair of points. Let $\mathcal{G}(\mathcal{V},\bm{W})$ correspond to a graph that consists of vertices $\mathcal{V}=\{v_{1},\ldots,v_{N}\}$ and edge weights $\bm{W}\in\mathbb{R}^{N\times N}$ such that $w_{ij}$ corresponds to some nonnegative weight, or similarity, between $v_{i}$ and $v_{j}$. The graph $\mathcal{G}$ is assumed to be undirected, i.e., $\bm{W}=\bm{W}^{T}$. The degree of a node is defined as $d_{i}=\sum_{j=1}^{N}w_{ij}$ and can be collected to form a degree matrix such that $\bm{D}=\text{diag}(d_{1},\ldots,d_{N})$. Let $\mathcal{A}_{1},\ldots,\mathcal{A}_{K}$ form a $K$-partition on $\mathcal{G}(\mathcal{V},\bm{W})$ and $\text{cut}(\mathcal{A},\mathcal{B})=\sum_{i\in\mathcal{A},j\in\mathcal{B}}w_{ij}$. Then, spectral clustering aims to solve the following optimization problem:
@@
where $\bar{\mathcal{A}}_{i}$ is all vertices not belonging in $\mathcal{A}_{i}$ and $|\mathcal{A}_{i}|$ is the number of vertices belonging to the $i$th partition. Instead of working with $\bm{W}$ directly, one computes a normalized graph Laplacian such as $\bm{L}=\bm{I}-\bm{D}^{-1}\bm{W}$ to make the influence of heavy degree nodes more similar to low degree nodes. Collecting the indicator values
$h_{ij}\triangleq\left\{\begin{array}[]{ll}1/\sqrt{|\mathcal{A}_{j}|}&\text{ if}v_{i}\in\mathcal{A}_{j}\\ 0&\text{ otherwise}\end{array}\right\}$
into a matrix $\bm{H}\in\mathbb{R}^{N\times K}$, one can rewrite (4) as
$\frac{1}{2}\sum_{i=1}^{K}\frac{\text{cut}(\mathcal{A}_{i},\bar{\mathcal{A}}_{i})}{|\mathcal{A}_{i}|}=\sum_{i=1}^{K}(\bm{H}^{T}\bm{L}\bm{H})_{ii}=\text{Tr}(\bm{H}^{T}\bm{L}\bm{H}),$ (5)
where $\text{Tr}(\bm{H})=\sum_{i}\bm{H}_{ii}$ denotes the trace of a matrix.
The indicator vectors are discrete, which makes the problem computationally challenging. One way to relax the problem is to allow arbitrary real values for $\bm{H}$ and instead solve
$\hat{\bm{H}}=\operatorname*{arg\,min}_{\bm{H}\in^{N\times K}}\text{Tr}(\bm{H}^{T}\bm{L}\bm{H})\text{ s.t. }\bm{H}^{T}\bm{H}=\bm{I}$ (6)
where the notation s.t. means subject to some conditions. Trace minimization problems with semi-unitary constraints are well-studied in the literature, and it is easy to see that $\hat{\bm{H}}$ consists of the first $K$ eigenvectors of $\bm{L}$ assuming that $\lambda_{1}(\bm{L})\leq\ldots\leq\lambda_{N}(\bm{L})$ where $\lambda_{i}(\bm{L})$ denotes the $i$th eigenvalue of $\bm{L}$. Once $\hat{\bm{H}}$ is computed, the $K$-means algorithm is applied to $\hat{\bm{H}}^{T}$, treating each row of $\hat{\bm{H}}$ as the spectral embedding of the associated $v_{i}$ vertex. The next section, Sec. III-A2, now describes some subspace clustering methods that use this technique by first forming a weight matrix $\bm{W}$ and then applying the spectral clustering method.
@@ III-A2 Self-Expressive Methods
Self-expressive methods exploit the “self-expressiveness property” [23] of data that hypothesizes that a single data point can be expressed as a linear combination of other data points in its cluster, which is trivially true if the data exactly follows a subspace model. The goal is to learn those linear coefficients and which is often achieved by adopting different regularizers in the formulation. These self-expressive algorithms, e.g., SSC [13] and LRSC [24], learn the coefficient matrix $\bm{P}\in\mathbb{R}^{N\times N}$ by solving special cases of the following general optimization problem
$\operatorname*{arg\,min}_{\bm{P}}f(\bm{Y}-\bm{Y}\bm{P})+\lambda\mathcal{R}(\bm{P})\hskip 5.69054pt\text{s.t.}\hskip 5.69054pt\bm{P}\in\mathcal{S}_{\bm{P}}.$ (7)
The function $f(\bm{Y}-\bm{Y}\bm{P})$ is a data fidelity term, $\mathcal{R}(\bm{P})$ is a regularizer, and $\mathcal{S}_{\bm{P}}$ is some constrained set to encourage $\bm{P}$ to satisfy certain conditions. In the case of SSC the optimization problem is
@@
where $\|\cdot\|_{1,1}=\|\text{vec}(\cdot)\|_{1}$ is the vectorized 1-norm and $\left\|\cdot\right\|_{\mathrm{F}}$ is the Frobenius norm of a matrix. Observe that (8) approximates each data sample as a sparse linear combination of other data points to form $\bm{P}$. Let $\text{abs}(\bm{P})$ represent the element-wise absolute value of matrix $\bm{P}$. Then, spectral clustering is performed on $\bm{W}=\frac{1}{2}(\text{abs}(\bm{P}^{T})+\text{abs}(\bm{P}))$ by applying the $K$-means method to the spectral embedding of the affinity matrix $\bm{W}$. Ideally, (8) would select the high quality data to represent worse samples in $\bm{P}$ and this information would be retained in $\bm{W}$. However, this data quality awareness condition is not guaranteed in self-expressive methods to our knowledge. In our experiments, Fig. 2(f) and Fig. 6 show that there must be an issue constructing an ideal $\bm{P}$ due to worse clustering performance in heteroscedastic conditions.

### v3

@@ Abstract
Principal component analysis (PCA) is a key tool in the field of data dimensionality reduction. Various methods have been proposed to extend PCA to the union of subspace (UoS) setting for clustering data that comes from multiple subspaces like $K$-Subspaces (KSS). However, some applications involve heterogeneous data that vary in quality due to noise characteristics associated with each data sample. Heteroscedastic methods aim to deal with such mixed data quality. This paper develops a heteroscedastic-based subspace clustering method, named ALPCAHUS, that can estimate the sample-wise noise variances and use this information to improve the estimate of the subspace bases associated with the low-rank structure of the data. This clustering algorithm builds on K-Subspaces (KSS) principles by extending the recently proposed heteroscedastic PCA method, named LR-ALPCAH, for clusters with heteroscedastic noise in the UoS setting. Simulations and real-data experiments show the effectiveness of accounting for data heteroscedasticity compared to existing clustering algorithms. Code available at https://github.com/javiersc1/ALPCAHUS.
@@ I Introduction
Many modern data science problems require learning an approximate signal subspace basis for some collection of data. This is important for downstream tasks involving subspace basis coefficients such as classification [classification], regression [regression], and compression [compression]. Besides subspace learning, one may be interested in clustering data points that originate from multiple subspaces. Formally, subspace clustering, or union of subspace (UoS) modeling, is an unsupervised machine learning problem where the goal is to cluster unlabeled data and find the subspaces associated with each data cluster. When the cluster assignments are known, it is easy to find the subspaces, and vice versa. This problem becomes nontrivial when both components must be estimated [subspace_review]. This clustering problem has many applications, such as image segmentation [image_segmentation], motion segmentation [motion_segmentation], image compression [image_compression], and system identification [systems_identification].
@@
Some applications involve heterogeneous data samples that vary in quality due in part to noise characteristics associated with each sample. A few examples of heteroscedastic datasets include environmental air quality data [epa], astronomical spectral data [astronomical_data], and biological sequencing data [bacteria_data]. In heteroscedastic settings, the noisier data samples can significantly corrupt the basis estimates [asymptotic_pca]. In turn, this corruption can worsen clustering performance as seen in Fig. 1. Popular clustering methods such as Sparse Subspace Clustering (SSC) [ssc], $K$-Subspaces (KSS) [k-subspaces], and Subspace Clustering via Thresholding (TSC) [heckel2015robust], all implicitly assume that data quality is consistent. For example, in SSC, the method relies on the self-expressiveness property of data that uses other similar samples to estimate each single sample. From our experiments, we found that this implicit data quality assumption can degrade clustering quality for heteroscedastic data.
@@
Figure 1: Two 1D subspaces, colored blue and yellow, with data consisting of two noise groups shown with circle (low noise) and triangle (high noise) markers.
@@
Because of these limitations, we developed a subspace clustering algorithm, inspired from KSS principles, that explicitly models noise variance terms, without assuming that data quality is known. The method adaptively clusters data while learning noise characteristics. See Fig. 1 for a visualization where Ensemble KSS (EKSS) [ekss] returns poor subspace bases estimates whereas our method found more accurate subspace bases and improved clustering quality.
@@
We extend our previous work [alpcah_journal] by generalizing the LR-ALPCAH formulation to the UoS setting for clustering heteroscedastic data. The proposed approach achieved ${\sim}3$ times lower clustering error than existing methods, and it achieved a relatively low clustering error even when very few high quality samples were available. The paper is divided into a few key sections. Sec. II introduces the heteroscedastic problem formulation for subspace clustering. Sec. III discusses related work in subspace clustering and reviews the heteroscedastic subspace algorithm LR-ALPCAH. Sec. IV introduces the proposed subspace clustering method named ALPCAHUS. Sec. V covers synthetic and real data experiments that illustrate the effectiveness of modeling heteroscedasticity in a clustering context. Finally, Sec. VI discusses some limitations of our method and possible extensions.
@@ II Problem Formulation
Let $K$ denote the number of subspaces that is either known beforehand or estimated using other methods. Before describing the general union of subspace model, the single subspace model under $K=1$ is introduced.
II-A Single Subspace Model ($K=1$)
Let $\bm{y}_{i}\in\mathbb{R}^{D}$ denote the data samples for index $i\in\{1,\ldots,N\}$, where $D$ denotes the ambient dimension and $N$ is the total number of samples. Let $\bm{x}_{i}$ represent the low-dimensional data sample generated by $\bm{x}_{i}=\bm{U}\bm{z}_{i}$ where $\bm{U}\in\mathbb{R}^{D\times d}$ is an unknown basis for a subspace of dimension $d$ and $\bm{z}_{i}\in\mathbb{R}^{d}$ are the corresponding basis coordinates. Then the heteroscedastic model we consider is
@@ II-A Single Subspace Model ($K=1$)
assuming Gaussian noise with variance $\nu_{i}$, where $\bm{I}$ denotes the $D\times D$ identity matrix. In this work, we consider the case where each data sample may have its own noise variance, however, one can adapt this method to consider the case where there are $G$ groups of data having shared noise variance terms $\{\nu_{1},\ldots,\nu_{G}\}$. Sec. III-C discusses an optimization problem based on this model that estimates the heterogeneous noise variances $\{\nu_{i}\}_{i=1}^{N}$ and the subspace basis $\bm{U}$.
@@ II-B Union of Subspaces Model ($K\geq 1$)
Let $\bm{Y}=\left[\begin{matrix}\bm{y}_{1}&\ldots&\bm{y}_{N}\end{matrix}\right]\in\mathbb{R}^{D\times N}$ denote a matrix whose columns consist of all $N$ data points $\bm{y}_{i}\in\mathbb{R}^{D}$. We generalize (1) to model the data with a union of subspaces model by
@@
where $\bm{U}_{k}\in\mathbb{R}^{D\times d_{k}}$ is a subspace basis that has subspace dimension $d_{k}$. Here, $\bm{z}_{i}\in\mathbb{R}^{d_{k}}$ denotes the basis coefficients associated with $\bm{x}_{i}$, and $\bm{\epsilon}_{i}\in\mathbb{R}^{D}$ denotes noise for that point drawn from $\bm{\epsilon}_{i}\sim\mathcal{N}(\bm{0},\nu_{i}\bm{I})$.
@@
$c_{i}=\operatorname*{arg\,min}_{k}\|\bm{y}_{i}-\bm{U}_{k}\bm{U}_{k}^{\prime}\bm{y}_{i}\|_{2}^{2},\quad\forall\bm{y}_{i}\in\bm{Y}.$ (3)
This label describes the subspace association of a data point $\bm{y}_{i}$ that has the lowest residual between the original sample and its reconstructed value given the subspace. In general, the goal is to estimate all of the subspace bases $\mathcal{U}=(\bm{U}_{1},\ldots,\bm{U}_{K})$ and cluster assignments $\mathcal{C}=(c_{1},\ldots,c_{N})$ associated with all the data samples.
@@ III-A Subspace Clustering
Many subspace clustering algorithms fall into a general umbrella of categories such as algebraic methods [algebraic_method], iterative methods [iterative_method], statistical methods [statistical_method], and spectral clustering methods [spectral_method1] [spectral_method2]. In recent years, both spectral clustering-based methods and iterative methods have become popular.
@@ III-A1 Spectral Clustering
Many methods build on spectral clustering. This method is often used for clustering nodes in graphs. Spectral clustering aims to find the minimum cost “cuts” in the graph to partition nodes into clusters. Given some collection of data points, one way to construct a graph is to assume that “nearby” data samples are highly connected in the graph. Thus, it is possible to construct a graph from data samples, with various levels of connectedness, by using some metric to assign affinity/similarity for each pair of points. Let $\mathcal{G}(\mathcal{V},\bm{W})$ correspond to a graph that consists of vertices $\mathcal{V}=\{v_{1},\ldots,v_{N}\}$ and edge weights $\bm{W}\in\mathbb{R}^{N\times N}$ such that $w_{ij}$ corresponds to some nonnegative weight, or similarity, between $v_{i}$ and $v_{j}$. The graph $\mathcal{G}$ is assumed to be undirected, i.e., $\bm{W}=\bm{W}^{\prime}$. The degree of a node is defined as $d_{i}=\sum_{j=1}^{N}w_{ij}$ and can be collected to form a degree matrix such that $\bm{D}=\text{diag}(d_{1},\ldots,d_{N})$. Let $\mathcal{A}_{1},\ldots,\mathcal{A}_{K}$ form a $K$-partition on $\mathcal{G}(\mathcal{V},\bm{W})$ and $\text{cut}(\mathcal{A},\mathcal{B})=\sum_{i\in\mathcal{A},j\in\mathcal{B}}w_{ij}$. Then, spectral clustering aims to solve the following
@@
where $\bar{\mathcal{A}}_{i}$ is all vertices not belonging in $\mathcal{A}_{i}$ and $|\mathcal{A}_{i}|$ is the number of vertices belonging to the $i$th partition. Instead of working with $\bm{W}$ directly, one computes a normalized graph Laplacian such as $\bm{L}=\bm{I}-\bm{D}^{-1}\bm{W}$ to make the influence of heavy degree nodes more similar to low degree nodes. Let $h_{i,j}$ indicate the $j$th cluster membership for the $i$th point by
$h_{ij}=\left\{\begin{array}[]{ll}1/\sqrt{|\mathcal{A}_{j}|}&\text{ if }v_{i}\in\mathcal{A}_{j}\\ 0&\text{ otherwise.}\end{array}\right$. (5)
Collecting the indicator values into a matrix $\bm{H}\in\mathbb{R}^{N\times K}$, one can rewrite (4) in matrix notation as
$\frac{1}{2}\sum_{i=1}^{K}\frac{\text{cut}(\mathcal{A}_{i},\bar{\mathcal{A}}_{i})}{|\mathcal{A}_{i}|}=\sum_{i=1}^{K}(\bm{H}^{\prime}\bm{L}\bm{H})_{ii}=\text{Tr}(\bm{H}^{\prime}\bm{L}\bm{H}),$ (6)
where $\text{Tr}(\bm{H})=\sum_{i}\bm{H}_{ii}$ denotes the trace of a matrix. The indicator vectors are discrete, which makes the problem computationally challenging. One way to relax the problem is to allow arbitrary real values for $\bm{H}$ and instead solve
$\hat{\bm{H}}=\operatorname*{arg\,min}_{\bm{H}\in^{N\times K}}\text{Tr}(\bm{H}^{\prime}\bm{L}\bm{H})\text{ s.t. }\bm{H}^{\prime}\bm{H}=\bm{I}$ (7)
where the notation s.t. means subject to some conditions. Trace minimization problems with semi-unitary constraints are well-studied in the literature, and it is easy to see that $\hat{\bm{H}}$ consists of the first $K$ eigenvectors of $\bm{L}$ associated with the smallest eigenvalues, i.e., $\lambda_{1}(\bm{L})\leq\ldots\leq\lambda_{N}(\bm{L})$ where $\lambda_{i}(\bm{L})$ denotes the $i$th eigenvalue of $\bm{L}$. Once $\hat{\bm{H}}$ is computed, the $K$-means algorithm is applied to $\hat{\bm{H}}^{\prime}$, treating each row of $\hat{\bm{H}}$ as the spectral embedding of the associated $v_{i}$ vertex. The next section, Sec. III-A2, now describes some subspace clustering methods that use this technique by first forming a weight matrix $\bm{W}$ and then applying the spectral clustering method on $\bm{W}$.
@@ III-A2 Self-Expressive Methods
Self-expressive methods exploit the “self-expressiveness property” [sc_survey] of data that hypothesizes that a single data point can be expressed as a linear combination of other data points in its cluster, which is trivially true if the data exactly follows a subspace model. The goal is to learn those linear coefficients which is often achieved by adopting different regularizers in the formulation. These self-expressive algorithms, e.g., SSC [ssc] and LRSC [lrsc], learn the coefficient matrix $\bm{P}\in\mathbb{R}^{N\times N}$ by solving special cases of the following general optimization problem
$\operatorname*{arg\,min}_{\bm{P}}\text{DF}(\bm{Y}-\bm{Y}\bm{P})+\lambda\Lambda(\bm{P})\hskip 5.69054pt\text{s.t.}\hskip 5.69054pt\bm{P}\in\mathcal{S}_{\bm{P}}.$ (8)
The function $\text{DF}(\bm{Y}-\bm{Y}\bm{P})$ is a data fidelity term, $\Lambda(\bm{P})$ is a regularizer, and $\mathcal{S}_{\bm{P}}$ is some constrained set to encourage $\bm{P}$ to satisfy certain conditions. In the case of SSC, the optimization problem is formulated as
@@
where $\|\bm{P}\|_{1,1}=\|\text{vec}(\bm{P})\|_{1}=\sum_{i}|p_{i}|$ is the vectorized 1-norm and $\left\|\cdot\right\|_{\mathrm{F}}=\|\text{vec}(\cdot)\|_{2}$ is the Frobenius norm of a matrix. Observe that (9) approximates each data sample as a sparse linear combination of other data points to form $\bm{P}$. Let $\text{abs}(\bm{P})$ represent the element-wise absolute value of matrix $\bm{P}$. Then, spectral clustering is performed on $\bm{W}=\frac{1}{2}(\text{abs}(\bm{P}^{\prime})+\text{abs}(\bm{P}))$ by applying the $K$-means method to the spectral embedding of the affinity matrix $\bm{W}$. Ideally, (9) would select the high quality data to represent worse samples in $\bm{P}$ and this information would be retained in $\bm{W}$. However, this data quality awareness condition is not guaranteed in self-expressive methods to our knowledge. In our experiments, Fig. 2(f) and Fig. 6 show that there must be an issue constructing an ideal $\bm{P}$ due to worse clustering performance in heteroscedastic conditions.